Graph the function using a ​[​-5,5,-​5,5] window and use the graph to estimate the domain and range.
​r(x)=|x|-5
1
Expert's answer
2020-11-24T08:16:28-0500




In its simplest form the domain is all the values that go into a function, and the range is all the values that come out.

Hence from the graph, we see domain is all real numbers, that is

"(-\infty ,\infty)"

and range is all values above -5 including -5.Hence range is "[-5,\infty )"